OpenStudy (anonymous):

I did an assignment about two weeks ago and my teacher gave me feedback & said I could resubmit. Can someone help me with what I did wrong? I'm not understanding this.

OpenStudy (anonymous):

@e.mccormick can you help please?

OpenStudy (e.mccormick):

Basically you would need to find where the trajectory and the orbit intersected. Substitution is a way of doing that. Hmm. Lets see.

OpenStudy (e.mccormick):

\(y^2 + x^2 = 40,000\) is the orbit You have \(y=4x-2\) as your trajectory. Since you know what y is equal to in the trajectory, you can repliace y in the orbit with that, then solve for x. Once you have the x you can put that in either equation and solve for y.

OpenStudy (anonymous):

Ohh, ohkay. I see. So I was missing a step. Thanks so much!

OpenStudy (e.mccormick):

No, it is a lot furthir out than that. \(x^2+y^2=40000\) is a circle with a radius \(r^2=40000\) so \(r=200\). That means the the numbers invold will have something pretty large in there, like (200,0) and (0,200) being on the circle.

OpenStudy (anonymous):

OH. Ohkay. I see now. I thought I was doing this wrong since it wasn't making since. Thank you.

OpenStudy (anonymous):

I have another quiz that I don't really know how to do, could you check my answers for me please?

OpenStudy (e.mccormick):

Here is a way you can check your answer when you try again: https://www.desmos.com/calculator/nkdqn63vwr If you hover the intersects it will give you an answer. You just need to work out the math to do it.

OpenStudy (anonymous):

That's another thing, I don't think my calculator was working. Thank you- thank you.

OpenStudy (e.mccormick):

You can use Google as a calculator. If you have an Android phone, RealCalc is a free scientific calculator ap. And there are several others on the web. And you said quiz? On the other, it was something already graded, so no big deal. But for non-graded quizzes and tests, you need to do your own work. OpenStudy can be used for homework or tutoring. Not quizzes and tests. But getting understanding of a graded quiz or test is fine because that is just tutoring.

OpenStudy (anonymous):

Ohkay, thank you! & yeah it's a practice quiz. It's Algebra 1 review I think. I just wanted to see if my answers were right because I'm not sure if I'm doing this right.

OpenStudy (e.mccormick):

Do you know what the change of base formula for logs is?

OpenStudy (anonymous):

I can't figure out how to draw on here, so:

OpenStudy (e.mccormick):

You can use \(\LaTeX\) for these: \(\log_ax=\dfrac{\log_bx}{\log_ba}\) So for solving something with a log like what you had with the interest, you can use change of base to put it into the calculator.

OpenStudy (anonymous):

Ohkay awesome. Thanks so much for all of your help.